What do you mean by projectile motion?

Grade:11

1 Answers

shashank Saxena
15 years ago

Projectile Motion:

A projectile is an object upon which the only force acting is gravity e.g. an object dropped from rest is a projectile (provided that the influence of air resistance is negligible), an object which is thrown vertically upward is also a projectile (provided that the influence of air resistance is negligible).


 



A projectile motion problem consists of two independent parts:

A.)Vertical motion at a constant downward acceleration, which is equal to a = -g = -9.80 m/s2

b.)Horizontal motion at a constant horizontal speed, vx = constant

Equations:
Vertical Motion: a= -g
                           Vy=v0sinθ (where v0 is the total velocity)
                           y-y0 = -1/2gt2 +v0y t
                           y = (vy2 - voy2)/(-2g)
                           vy - voy = -gt
                          
Horizontal Motion: a = 0
                               Vx =v0cosθ          
                                x - x0 = vxt

Frame of reference:

Projectile motion problems need a two dimensional coordinate system to describe the projectile's motion. The y-direction is usually associated with the vertical motion and location of the projectile, while the x-direction is usually associated with the horizontal motion and location of the projectile.* The generic equations for projectile motion assume the origin is at ground level, up is positive, right is positive, and that the clock starts the moment the projectile leaves the ground